OpenStudy (anonymous):

What is the slope of the graph of 3x + 6y = 36?

OpenStudy (madhu.mukherjee.946):

6y=36-3x y=6-x/2 so ,your slope is -1/2

OpenStudy (anonymous):

thank you again madam :)

OpenStudy (madhu.mukherjee.946):

your welcome:)